# Language Models as Hierarchy Encoders

### v1

@@ Abstract
Interpreting hierarchical structures latent in language is a key limitation of current language models (LMs). While previous research has implicitly leveraged these hierarchies to enhance LMs, approaches for their explicit encoding are yet to be explored. To address this, we introduce a novel approach to re-train transformer encoder-based LMs as Hierarchy Transformer encoders (HiTs), harnessing the expansive nature of hyperbolic space. Our method situates the output embedding space of pre-trained LMs within a Poincaré ball with a curvature that adapts to the embedding dimension, followed by re-training on hyperbolic cluster and centripetal losses. These losses are designed to effectively cluster related entities (input as texts) and organise them hierarchically. We evaluate HiTs against pre-trained and fine-tuned LMs, focusing on their capabilities in simulating transitive inference, predicting subsumptions, and transferring knowledge across hierarchies. The results demonstrate that HiTs consistently outperform both pre-trained and fine-tuned LMs in these tasks, underscoring the effectiveness and transferability of our re-trained hierarchy encoders.222 Code on GitHub: https://github.com/KRR-Oxford/HierarchyTransformers; Datasets on Zenodo: https://zenodo.org/doi/10.5281/zenodo.10511042; HiT models on Huggingface Hub: https://huggingface.co/Hierarchy-Transformers
Language Models, Natural Language Processing, Deep Learning, Machine Learning, ICML
@@ 1 Introduction
Figure 1: Illustration of how hierarchies are explicitly encoded in HiT after re-training. The square ($d$-dimensional hyper-cube) refers to the output embedding space of transformer encoder-based LMs whose final activation function is $\tanh$, and the circumscribed circle ($d$-dimensional hyper-sphere) refers to the Poincaré ball of radius $\sqrt{d}$. The distance and norm metrics involved in our re-training hyperbolic losses are defined w.r.t. this manifold.
@@
In the field of Natural Language Processing (NLP) and related areas, the emergence of transformer-based language models (LMs) such as BERT (encoder-based) (Kenton & Toutanova, 2019), GPT (decoder-based) (Radford et al., 2018), and the more recent large language models (LLMs) like GPT-4 (OpenAI, 2023) and Llama 2 (Touvron et al., 2023), has marked a significant progression. Nonetheless, these models face a notable challenge in effectively encoding and interpreting hierarchical structures latent in language. This limitation has been highlighted by several studies, including those by Hanna & Mareček (2021) and He et al. (2023b), which employed prompt-based probes to reveal the limited hierarchical knowledge in pre-trained LMs, and the work by Lin & Ng (2022), which demonstrated these models’ struggles with capturing the transitivity of hierarchical relationships.
@@
Prior research has explored various methods to infuse hierarchical information into LM training. Common approaches include classification-based fine-tuning using sentence head embedding with a classification layer (Chen et al., 2023) or few-shot prompting with an answer mapping to classification labels (He et al., 2023b). To pre-train, or re-train [^3] The term re-train refers to further pre-train LMs on another corpus; it is distinguished from fine-tune as it does not modify the LM’s architecture, while the latter often involves adding task-specific layers which lead to additional learnable parameters. LMs on a corpus constructed from hierarchical data, Liu et al. (2020) converted structural representations into textual formats to align the masked language modeling objective. Others, like Liu et al. (2021) and Gosselin & Zouaq (2023), have focused on extracting analogous and contrasting examples from hierarchical structures for a similarity-based contrastive learning objective. The aforementioned studies leveraged hierarchical information as implicit signals to augment LMs, yet no existing works specifically targeted the explicit encoding of hierarchies with LMs.
@@
To bridge this gap, we introduce a novel approach to re-train transformer encoder-based LMs as Hierarchy Transformer encoders (HiTs). Inspired by the efficacy of hyperbolic geometry in representing hierarchical structures (Nickel & Kiela, 2017; Ganea et al., 2018b), we propose the hyperbolic clustering and centripetal losses tailored for LM re-training. As illustrated in Figure 1, transformer encoder-based LMs typically use a $\tanh$ activation function in the last layer, which maps each embedding dimension to the range $[-1,1]$. Consequently, the output embeddings of LMs are confined within a unit $d$-dimensional hyper-cube. Leveraging this characteristic, we utilise a Poincaré ball of radius $\sqrt{d}$, whose boundary circumscribes [^4] We ignore the vertices of the hyper-cube, as they are on the boundary and thus are undefined in the open Poincaré ball. the output embedding space of LMs. The metrics for distance and norm used in our hyperbolic losses are defined w.r.t. this specific manifold. After re-training, entities are not only clustered according to their relatedness but also hierarchically organised.
@@
In evaluating HiTs, we compare their performance against both pre-trained and fine-tuned LMs in the Multi-hop Inference and Mixed-hop Prediction tasks. The Multi-hop Inference task, following the setting in Ganea et al. (2018b), involves training models on all asserted (i.e., one-hop) subsumptions and assessing their ability to infer transitive (i.e., multi-hop) subsumptions. The Mixed-hop Prediction task is designed to mirror real-world scenarios, where models trained on incomplete hierarchy are applied to predict unknown subsumption relationships between arbitrary entity pairs. Additionally, we introduce a transfer learning setting, where models trained on one hierarchy are tested on another. Our experiments utilise datasets derived from WordNet’s noun hierarchy (Miller, 1995) and the SNOMED CT ontology (Stearns et al., 2001), [^5] Results on WordNet (Noun) are presented in Section 4, while results on SNOMED CT are presented in Appendix C. and transfer evaluation datasets from Schema.org (Guha et al., 2016), Food Ontology (FoodOn) (Dooley et al., 2018), and Disease Ontology (DOID) (Schriml et al., 2012). The results show that HiTs significantly surpass both pre-trained and fine-tuned LMs in these tasks, demonstrating their robustness to generalise from asserted to inferred and unseen subsumptions, and a promising potential in hierarchy-oriented semantic search.
@@ 2 Preliminaries
Before delving into our methodology, we provide a brief overview of key concepts: transformer encoder-based LMs, hyperbolic geometry, and the formal definition of hierarchy.
@@ 2.1 Language Models
Transformer encoder-based LMs excel in providing fine-grained contextual word embeddings for enhanced language understanding. A key component of these models is the self-attention mechanism, which dynamically assigns importance to different segments of the input text, thereby capturing nuanced contextual semantics more effectively. Notable examples of such models include BERT (Kenton & Toutanova, 2019) and RoBERTa (Liu et al., 2019), both of which utilise the masked language modelling objective during pre-training. This approach involves partially masking input sentences and prompting the model to predict the masked tokens, using the unmasked surrounding text as context. For acquiring sentence-level embeddings, these models can be augmented with an additional pooling layer, applied over the token embeddings (Xiao, 2018; Reimers & Gurevych, 2019). Pooling strategies such as mean, max, and sentence head pooling are employed, with their effectiveness varying across different applications. A contrastive learning objective is often applied for refining sentence-level semantics (Reimers & Gurevych, 2019; Gao et al., 2021). Despite the rise of generative LLMs, transformer encoder-based LMs maintain their importance, offering versatility and efficiency in tasks like text classification and semantic search.
@@ 2.2 Hyperbolic Geometry
Hyperbolic geometry, a form of non-Euclidean geometry, is characterised by its constant negative Gaussian curvature, a fundamental aspect that differentiates it from the flat, zero curvature of Euclidean geometry. In hyperbolic space, distances between points increase exponentially as one moves towards the boundary, making it inherently suitable for embedding hierarchical structures. This intuition aligns with the tree embedding theorem based on $\delta$-hyperbolicity, as discussed in Gromov (1987) and Ganea et al. (2018b).
@@ 3 Hierarchy Transformer Encoder
We intend to propose a general and effective strategy to re-train transformer encoder-based LMs as Hierarchy Transformer encoders (HiTs). To deal with arbitrary input lengths of entity names, we employ an architecture similar to sentence transformers (Reimers & Gurevych, 2019), incorporating a mean pooling layer over token embeddings to produce sentence embeddings for entities. Note that some of the sentence transformer models have a normalisation layer after pooling; we exclude this layer because its presence will constrain the embeddings’ Euclidean norms to one, thus hindering hierarchical organisation of entity embeddings. It is also worth mentioning that these changes do not add in learnable parameters besides the ones already in LMs.
As aforementioned, the output embedding space of these LMs is typically a $d$-dimensional hyper-cube because of the $\tanh$ activation function in the last layer. Thus, we can construct a Poincaré ball of radius $\sqrt{d}$ (or equivalently, curvature value $c=\frac{1}{d}$) whose boundary circumscribes the hyper-cube. [^8] We have also considered scaling down each dimension of the LM embeddings by $\sqrt{d}$ to confine them within a unit Poincaré ball, but we found that losses are harder to converge in this construction. Based on this manifold, we propose the following two losses for hierarchy re-training.
Figure 2: Illustration of the impact of hyperbolic clustering and centripetal losses during training. In Euclidean space, it seems contradictory that both “phone” and “computer” are pulled towards “e-device” but are also pushed away from each other. However, in principle this is not a problem in hyperbolic space, where distances increase exponentially relative to Euclidean distances as one moves from the origin to the boundary of the manifold.
@@
Here, inputs are presented in the form of triplet $(e,e^{+},e^{-})$, where $e^{+}$ is a parent entity of $e$, and $e^{-}$ is a negative parent entity of $e$; $\mathcal{D}$ denotes the set of these triplets; $d_{c}(\cdot,\cdot)$ refers to the hyperbolic distance function defined in equation 1, and $\alpha$ is the hyperbolic distance margin. The bold letters denote the embeddings of the corresponding entities.
@@
In Figure 2, we demonstrate the impact of $\mathcal{L}_{\textsc{HiT}}$ on entity embeddings. The entity “e-device”, being most general, is nearest to the origin. Sibling entities, such as “phone” and “computer”, “laptop” and “pc”, are closer to their common parent than to each other, illustrating the effect of retraining to cluster related entities while maintain hierarchical relationships.
As the HiT model functions as an encoder, it does not inherently support direct predictions of subsumption relationships. To address this, we devise an empirical scoring function that leverages the hierarchy re-trained entity embeddings. This function aims to predict the subsumption relationship for any given pair of entities $(e_{1},e_{2})$, incorporating both the clustering and centripetal heuristics:
@@ 4.1 Task Definition
In our evaluation, we propose the following two tasks to assess a model’s ability to generalise from asserted to inferred and unseen subsumptions.
@@
This task, following the setting in Ganea et al. (2018b), aims to evaluate the model’s ability in deducing indirect, multi-hop subsumptions $\mathcal{T}$ from direct, one-hop subsumptions $\mathcal{E}$, so as to simulate transitive inference. We split $\mathcal{T}$ for validation and testing, denoted as $\mathcal{T}_{val}$ and $\mathcal{T}_{test}$, respectively. For each positive subsumption $(e,e^{+})$ involved, we sampled $10$ negative parents $e^{-}$ for $e$, leading to $10$ training triplets99910101010 training triplets are constructed from $11$ entity pairs.. Following the criteria in Section 2.3, $(e,e^{-})$ is a valid negative if $(e,e^{-})\not\in\mathcal{E}\cup\mathcal{T}$. We further split the task into two settings: one with random negatives and another with hard negatives, the latter mainly comprising sibling entities. Since not every entity has enough siblings, we supplemented with random negatives that have been sampled to maintain a consistent positive-to-negative ratio of $1:10$.
@@
This task aims to evaluate the model’s capability in determining the existence of subsumption relationships between arbitrary entity pairs, where the entities are not necessarily seen during training. We propose a challenging setting where models are trained on incomplete direct subsumptions and examined on a mix of hold-out, unseen direct and indirect (mixed-hop) subsumptions. We split $\mathcal{E}$ into training, validation, and test sets, denoted as $\mathcal{E}_{train}$, $\mathcal{E}_{val}$, and $\mathcal{E}_{test}$, respectively. The final training, validataion, and test sets for this task are $\mathcal{E}_{train}$, $\mathcal{E}_{val}\cup\mathcal{T}_{val}$, and $\mathcal{E}_{test}\cup\mathcal{T}_{test}$, respectively, where $\mathcal{T}_{val}$ and $\mathcal{T}_{test}$ are re-used from the previous task. Again, each positive subsumption in these sets is paired with $10$ negative samples, either randomly chosen or from sibling entities. Furthermore, an important factor that reflects the model’s generalisability is to examine the transfer ability across hierarchies. To this end, we extend the mixed-hop prediction task with a transfer setting where models trained on one hierarchy are tested on another.
@@ 4.2 Dataset Construction
We constructed the primary dataset from the noun hierarchy of WordNet (Miller, 1995) due to its comprehensive and structured representation of linguistic hierarchies. To assess the transferability and robustness across different domains, we additionally constructed datasets from ontologies that represent varied semantic granularities and domains, namely Schema.org (Guha et al., 2016), Food Ontology (FoodOn) (Dooley et al., 2018), and Disease Ontology (DOID) (Schriml et al., 2012). We retrieved WordNet from NLTK [^10] https://www.nltk.org/ and adopted pre-processing steps similar to Nickel & Kiela (2017), utilising the hypernym relations between noun synsets to construct the hierarchy. For Schema.org, FoodOn, and DOID, our pre-processing paralleled that in He et al. (2023b), transforming these ontologies into hierarchies of named entities (details in Appendix A). To accommodate the textual input requirements of LMs, we constructed an entity lexicon using the name attribute in WordNet and the rdfs:label property in the ontologies. [^11] We selected the first name (in English) if multiple names for one entity were available.
@@
In addition to our main evaluation, we constructed a dataset from the widely-recognised biomedical ontology SNOMED CT (Stearns et al., 2001) and conducted futher evaluation. The relevant details are presented in Appendix C.

### v2

@@ Abstract
Interpreting hierarchical structures latent in language is a key limitation of current language models (LMs). While previous research has implicitly leveraged these hierarchies to enhance LMs, approaches for their explicit encoding are yet to be explored. To address this, we introduce a novel approach to re-train transformer encoder-based LMs as Hierarchy Transformer encoders (HiTs), harnessing the expansive nature of hyperbolic space. Our method situates the output embedding space of pre-trained LMs within a Poincaré ball with a curvature that adapts to the embedding dimension, followed by training on hyperbolic clustering and centripetal losses. These losses are designed to effectively cluster related entities (input as texts) and organise them hierarchically. We evaluate HiTs against pre-trained LMs, standard fine-tuned LMs, and several hyperbolic embedding baselines, focusing on their capabilities in simulating transitive inference, predicting subsumptions, and transferring knowledge across hierarchies. The results demonstrate that HiTs consistently outperform all baselines in these tasks, underscoring the effectiveness and transferability of our re-trained hierarchy encoders.111 See GitHub repository: https://github.com/KRR-Oxford/HierarchyTransformers; Datasets on Zenodo: https://zenodo.org/doi/10.5281/zenodo.10511042; and HiT models on the Huggingface Hub: https://huggingface.co/Hierarchy-Transformers.
@@ 1 Introduction
In the field of Natural Language Processing (NLP) and related areas, the emergence of transformer-based language models (LMs) such as BERT (encoder-based) [1], GPT (decoder-based) [2], and the more recent large language models (LLMs) like GPT-4 [3] and Llama 2 [4], has marked a significant progression. Nonetheless, these models face a notable challenge in effectively encoding and interpreting hierarchical structures latent in language. This limitation has been highlighted by several studies, including those by [5] and [6], which employed prompt-based probes to reveal the limited hierarchical knowledge in pre-trained LMs, and the work by [7], which demonstrated these models’ struggles with capturing the transitivity of hierarchical relationships.
@@
Prior research has explored various methods to infuse hierarchical information into LM training. Common approaches include classification-based fine-tuning using sentence head embedding with a classification layer [8] or few-shot prompting with an answer mapping to classification labels [6]. To further pre-train, or re-train [^2] In this work, the term re-train refers to train LMs on a new corpus without modifying its architecture; it is distinguished from standard fine-tuning that involves adding task-specific layers which lead to additional learnable parameters. LMs on a corpus constructed from hierarchical data, [9] converted structural representations into textual formats to align the masked language modeling objective. Others, like [10] and [11], have focused on extracting analogous and contrasting examples from hierarchical structures for a similarity-based contrastive learning objective. The aforementioned studies leveraged hierarchical information as implicit signals to augment LMs, yet no existing works specifically targeted the explicit encoding of hierarchies with LMs.
Figure 1: Illustration of how hierarchies are explicitly encoded in HiTs. The square ($d$-dimensional hyper-cube) refers to the output embedding space of transformer encoder-based LMs whose final activation function is typically $\tanh$, and the circumscribed circle ($d$-dimensional hyper-sphere) refers to the Poincaré ball of radius $\sqrt{d}$. The distance and norm metrics involved in our hyperbolic losses are defined w.r.t. this manifold.
@@
To bridge this gap, we introduce a novel approach to re-train transformer encoder-based LMs as Hierarchy Transformer encoders (HiTs). Inspired by the efficacy of hyperbolic geometry in representing hierarchical structures [12, 13], we propose the hyperbolic clustering and centripetal losses tailored for LM re-training. As illustrated in Figure 1, transformer encoder-based LMs typically use a $\tanh$ activation function in the last layer, which maps each embedding dimension to the range $[-1,1]$. Consequently, the output embeddings of LMs are confined within a unit $d$-dimensional hyper-cube. Leveraging this characteristic, we utilise a Poincaré ball of radius $\sqrt{d}$, whose boundary circumscribes [^3] We ignore the vertices of the hyper-cube, as they are on the boundary and thus are undefined in the open Poincaré ball. the output embedding space of LMs. The metrics for distance and norm used in our hyperbolic losses are defined w.r.t. this specific manifold. After re-training, entities are not only clustered according to their relatedness but also hierarchically organised.
@@
In evaluating HiTs, we compare their performance against pre-trained LMs, standard fine-tuned LMs, and previous hyperbolic embedding models in the Multi-hop Inference and Mixed-hop Prediction tasks. The Multi-hop Inference task, following the setting in [13], involves training models on all asserted (i.e., one-hop) subsumptions and assessing their ability to infer transitive (i.e., multi-hop) subsumptions. The Mixed-hop Prediction task is designed to mirror real-world scenarios, where models trained on incomplete hierarchy are applied to predict unknown subsumption relationships between arbitrary entity pairs. Additionally, we introduce a transfer learning setting, where models trained on one hierarchy are tested on another. Our experiments utilise datasets derived from WordNet [14] and SNOMED CT [15], [^4] Results on WordNet are presented in Section 4, while results on SNOMED CT are presented in Appendix D. and transfer evaluation datasets from Schema.org [16], Food Ontology (FoodOn) [17], and Disease Ontology (DOID) [18]. The results show that HiTs significantly surpass all baselines in these tasks, demonstrating their robustness to generalise from asserted to inferred and unseen subsumptions, and a promising potential in hierarchy-oriented semantic search.
@@ 2.1 Language Models
Transformer encoder-based LMs excel in providing fine-grained contextual word embeddings for enhanced language understanding. A key component of these models is the self-attention mechanism, which dynamically assigns importance to different segments of the input text, thereby capturing nuanced contextual semantics more effectively. Notable examples of such models include BERT [1] and RoBERTa [19], both of which utilise the masked language modelling objective during pre-training. This approach involves partially masking input sentences and prompting the model to predict the masked tokens, using the unmasked surrounding text as context. For acquiring sentence-level embeddings, these models can be augmented with an additional pooling layer, applied over the token embeddings [20, 21]. Pooling strategies such as mean, max, and sentence head pooling are employed, with their effectiveness varying across different applications. A contrastive learning objective is often applied for refining sentence-level semantics [21, 22]. Despite the rise of generative LLMs, transformer encoder-based LMs maintain their importance, offering versatility and efficiency in tasks like text classification and semantic search.
@@ 2.2 Hyperbolic Geometry
Hyperbolic geometry, a form of non-Euclidean geometry, is featured by its constant negative Gaussian curvature, a fundamental aspect that differentiates it from the flat, zero curvature of Euclidean geometry. In hyperbolic space, distances between points increase exponentially as one moves towards the boundary, making it inherently suitable for embedding hierarchical structures. This intuition aligns with the tree embedding theorem based on $\delta$-hyperbolicity, as discussed in [23] and [13].
@@ 3 Hierarchy Transformer Encoder
We intend to propose a general and effective strategy to re-train transformer encoder-based LMs as Hierarchy Transformer encoders (HiTs). To deal with arbitrary input lengths of entity names, we employ an architecture similar to sentence transformers [21], incorporating a mean pooling layer over token embeddings to produce sentence embeddings for entities. Note that some of the sentence transformer models have a normalisation layer after pooling; we exclude this layer because its presence will constrain the embeddings’ Euclidean norms to one, thus hindering hierarchical organisation of entity embeddings. It is also worth mentioning that these changes do not add in learnable parameters besides the ones already in LMs, thus retaining the original architectures of LMs as encoders. As aforementioned, the output embedding space of these LMs is typically a $d$-dimensional hyper-cube because of the $\tanh$ activation function in the last layer. Thus, we can construct a Poincaré ball of radius $\sqrt{d}$ (or equivalently, curvature value $c=\frac{1}{d}$) whose boundary circumscribes the hyper-cube. [^7] We have also considered scaling down each dimension of the LM embeddings by $\sqrt{d}$ to confine them within a unit Poincaré ball, but we found that losses are harder to converge in this construction. Unlike previous hyperbolic embedding methods that utilise the entire hyperbolic space and often require a projection layer to manage out-of-manifold embeddings, our method ensures that embeddings are contained within a specific subset of this manifold. Empirical evidence supports that this subset sufficiently accommodates entities in high-dimensional space (see Section 4.5). Based on this curvature-adapted manifold, we propose the following two losses for hierarchy re-training.
@@
Here, inputs are presented in the form of triplet $(e,e^{+},e^{-})$, where $e^{+}$ is a parent entity of $e$, and $e^{-}$ is a negative parent entity of $e$; $\mathcal{D}$ denotes the set of these triplets; $d_{c}(\cdot,\cdot)$ refers to the hyperbolic distance function defined in Equation (1), and $\alpha$ is the hyperbolic distance margin. The bold letters denote the embeddings of the corresponding entities.
@@
Figure 2: Illustration of the impact of $\mathcal{L}_{\textsc{HiT}}$ during training. In Euclidean space, it seems contradictory that both “phone” and “computer” are pulled towards “e-device” but are also pushed away from each other. However, in principle this is not a problem in hyperbolic space, where distances increase exponentially relative to Euclidean distances as one moves from the origin to the boundary of the manifold.
In Figure 2, we demonstrate the impact of $\mathcal{L}_{\textsc{HiT}}$ on entity embeddings. The entity “e-device”, being most general, is nearest to the origin. Sibling entities, such as “phone” and “computer”, “laptop” and “pc”, are closer to their common parent than to each other, illustrating the effect of re-training to cluster related entities while maintain hierarchical relationships.
As the HiT model functions as an encoder, it does not inherently support direct predictions of subsumption relationships. To address this, we devise a probing function that leverages the hierarchy re-trained entity embeddings. This function aims to predict the subsumption relationship for any given pair of entities $(e_{1},e_{2})$, incorporating both the clustering and centripetal heuristics:
@@ 4.1 Task Definition
This task, following the setting in [13], aims to evaluate the model’s ability in deducing indirect, multi-hop subsumptions $\mathcal{T}$ from direct, one-hop subsumptions $\mathcal{E}$, so as to simulate transitive inference. We split $\mathcal{T}$ for validation and testing, denoted as $\mathcal{T}_{val}$ and $\mathcal{T}_{test}$, respectively. For each positive subsumption $(e,e^{+})$ involved, we sampled $10$ negative parents $e^{-}$ for $e$, leading to $10$ training triplets88810101010 training triplets are constructed from $11$ entity pairs.. Following the criteria in Section 2.3, $(e,e^{-})$ is a valid negative if $(e,e^{-})\not\in\mathcal{E}\cup\mathcal{T}$. We further split the task into two settings: one with random negatives and another with hard negatives, the latter mainly comprising sibling entities. Since not every entity has enough siblings, we supplemented with random negatives that have been sampled to maintain a consistent positive-to-negative ratio of $1:10$.
@@
This task aims to evaluate the model’s capability in determining the existence of subsumption relationships between arbitrary entity pairs, where the entities are not necessarily seen during training. We propose a challenging setting where models are trained on incomplete direct subsumptions and examined on a mix of hold-out, unseen direct and indirect (mixed-hop) subsumptions. We split $\mathcal{E}$ into training, validation, and test sets, denoted as $\mathcal{E}_{train}$, $\mathcal{E}_{val}$, and $\mathcal{E}_{test}$, respectively. The final training, validataion, and test sets for this task are $\mathcal{E}_{train}$, $\mathcal{E}_{val}\cup\mathcal{T}_{val}$, and $\mathcal{E}_{test}\cup\mathcal{T}_{test}$, respectively, where $\mathcal{T}_{val}$ and $\mathcal{T}_{test}$ are re-used from the previous task. Again, each positive subsumption in these sets is paired with $10$ negative samples, either randomly chosen or from sibling entities. Furthermore, an important factor that reflects the model’s generalisability is to examine the transfer ability across hierarchies. To this end, we extend the mixed-hop prediction task with a transfer setting where models trained on asserted training edges of one hierarchy are tested on arbitrary entity pairs of another.
@@ 4.2 Dataset Construction
We constructed the primary dataset from the noun hierarchy of WordNet [14] due to its comprehensive and structured representation of linguistic hierarchies. To assess the transferability and robustness across different domains, we additionally constructed datasets from ontologies that represent varied semantic granularities and domains, namely Schema.org [16], Food Ontology (FoodOn) [17], and Disease Ontology (DOID) [18]. We retrieved WordNet from NLTK [25] and adopted pre-processing steps similar to [12], utilising the hypernym relations between noun synsets to construct the hierarchy. For Schema.org, FoodOn, and DOID, our pre-processing paralleled that in [6], transforming these ontologies into hierarchies of named entities (details in Appendix A). To accommodate the textual input requirements of LMs, we constructed an entity lexicon using the name attribute in WordNet and the rdfs:label property in the ontologies. [^9] We selected the first name (in English) if multiple names for one entity were available.
@@
In addition to our main evaluation, we constructed a dataset from the widely-recognised biomedical ontology SNOMED CT [15] and conducted futher evaluation. The relevant details are presented in Appendix D.
4.3 Baselines
Naive Prior
We first introduce a naive baseline (NaivePrior) that utilises the prior probability of positive subsumptions in the training set for prediction. Given that each positive sample is paired with $10$ negatives, the prior probability of a positive prediction stands at $\frac{1}{11}$. Consequently, Precision, Recall, and F-score on the test set are all $\frac{1}{11}$.
Pre-trained LMs
[…]
Fine-tuned LMs
Fine-tuned LMs are used as baselines to demonstrate that despite the efficacy in various tasks, standard fine-tuning struggles to address this specific challenge. Following the BERTSubs approach outlined in [8], we employ pre-trained LMs with an added linear layer for binary classification, and optimising on the Softmax loss. [8] have shown that this method outperforms various structure-based embeddings such as TransE [28] and DistMult [29], and also surpasses OWL2Vec* [30], which integrates both structural and textual embeddings, in subsumption prediction.
Hyperbolic Baselines
Previous static hyperbolic embedding models are typically evaluated using the Multi-hop Inference task. In our study, we select the Poincaré Embedding (PoincaréEmbed) [12] and the Hyperbolic Entailment Cone (HyperbolicCone) [13] as baselines on this task. However, their lack of inductive prediction capabilities prevents their evaluation on the Mixed-hop Prediction task and its transfer setting. Additionally, we include the hyperbolic GloVe embedding (PoincaréGloVe) as a baseline in our transfer evaluation. We select the best-performing PoincaréGloVe (50$\times$2D with an initial trick) pre-trained on a 1.4B token English Wikipedia dump. While PoincaréGloVe supports inductive prediction, its effectiveness is limited by word-level tokenisation, rendering it less effective at handling unknown words. To address this, we employ NaivePrior as a fallback method for entities that involve unknown words and cannot be predicted by PoincaréGloVe.
More details of our code implementation and experiment settings are presented in Appendix B.
Table 2: Multi-hop Inference and Mixed-hop Prediction test results on WordNet.